# A replica analysis of under-bagging

### v1

@@ Abstract
A sharp asymptotics of the under-bagging (UB) method, which is a popular ensemble learning method for training classifiers from an imbalanced data, is derived and used to compare with several other standard methods for learning from imbalanced data, in the scenario where a linear classifier is trained from a binary mixture data. The methods compared include the under-sampling (US) method, which trains a model using a single realization of the subsampled dataset, and the simple weighting (SW) method, which trains a model with a weighted loss on the entire data. It is shown that the performance of UB is improved by increasing the size of the majority class, even if the class imbalance can be large, especially when the size of the minority class is small. This is in contrast to US, whose performance does not change as the size of the majority class increases, and SW, whose performance decreases as the imbalance increases. These results are different from the case of the naive bagging in training generalized linear models without considering the structure of class imbalance, indicating the intrinsic difference between the ensembling and the direct regularization on the parameters.
@@ 1 Introduction
In the context of binary classification, class imbalance refers to the situations where one class overwhelms the other in terms of the number of data points. Learning on class imbalanced data often occurs in real-world classification tasks, which occur in diverse domains such as the medical data analysis (Cohen et al., 2006), the image classification tasks (Khan et al., 2018), and so on. The primary goal when training classifiers on such imbalanced data is to achieve good generalization to both minority and majority classes. However, standard supervised learning without any special care often results in poor generalization performance for the minority class. Therefore, special treatments are necessary and have been one of the major research topics in the machine learning (ML) community for decades.
@@
Under-bagging (UB) (Wallace et al., 2011) is a popular and efficient method for dealing with a class imbalance that combines under-sampling (US) and bagging (Breiman, 1996). The basic idea of UB is to address label imbalance by randomly discarding a portion of the majority class data, ensuring an equal number of data points between the minority and majority classes. However, this data reduction increases the variance of the estimator since the total number of training data is reduced. To mitigate this, UB aggregates estimates for different under-sampled instances, similar to the aggregation performed in bootstrap aggregation (Breiman, 1996). US corresponds to a method of training a classifier on a single undersampled data. (Wallace et al., 2011) argues that UB is particularly effective in the training of overparametrized models, where the number of the model parameters far exceeds the size of the training data.
@@
There are also another line of approach, that is, cost-sensitive methods, which modify the training loss to account for varying class penalties. However, it has been recently pointed out that training overparametrized models using cost-sensitive methods requires delicate design of loss functions (Lin et al., 2017; Khan et al., 2018; Byrd & Lipton, 2019; Ye et al., 2020; Kini et al., 2021) instead of the naive weighted class-entropy and so on. In contrast to such delicate cost-sensitive approaches, UB has the advantage of being relatively straightforward to use, as it achieves complete class balance in each under-sampled dataset.
@@
Bagging is a natural approach to reduce the increased variance of weak learners, resulting from the smaller data size due to the resampling. However, unlike decision trees, which are the primary models used in bagging, neural networks can use simple regularization methods such as ridge regularization to reduce the variance. In fact, is has been experimentally reported that naive bagging, which ignores the structure of class balance, does not improve the classification performance of deep neural networks (Nixon et al., 2020). Furthermore, in generalized linear models (GLMs), it has been both theoretically (Sollich & Krogh, 1995; Krogh & Sollich, 1997; LeJeune et al., 2020; Du et al., 2023) and experimentally (Yao et al., 2021) reported that bagging with a specific sampling rate is exactly equivalent to the ridge regularization with properly adjusted regularization parameter. These results indicate that naive bagging plays the role of an implicit regularization in linear models. Given these results, the question arises whether it is better to use UB, which requires an increased computational cost proportional to the number of under-sampled datasets, when training linear models or neural networks, rather than just employing ridge regularization.
@@
Taking into account the aforementioned situation, this study aims to clarify whether bagging or ridge regularization should be used when learning from class-imbalanced data. Specifically, we compare the performance of classifiers obtained by UB with those obtained by using a single under-sampled data with appropriate ridge regularization, which we refer to as the simple weighting (SW) method, when training a linear model on binary mixture data, through the $F$-measure. To do this, we derive a sharp characterization of the classifiers obtained by the randomly subsampled data in the asymptotic limit where the input dimension and the parameter dimension diverge at the same rate, and then use it to analyze the performance of UB, US and SW.
@@
  • A sharp characterization of the statistical behavior of the linear classifier is derived, in which the classifier is obtained by minimizing the reweighted empirical risk function on binary mixture data, in the asymptotic limit where input dimension and data size diverge proportionally (Claim 4). This includes the properties of the estimator obtained from the under-sampled data and the simple weighted loss function. There, statistical property of the logits for a new input is described by a small finite set of scalar quantities determined as a solution of nonlinear equations, which we refer to as self-consistent equations (Claim 1). The sharp characterization of $F$-measure is also obtained (Claim 5). The derivation is based on the replica method of statistical mechanics (Mézard et al., 1987; Charbonneau et al., 2023; Montanari & Sen, 2024).
  • By using the derived sharp asymptotics, it is shown that the estimators obtained by UB have higher generalization performance in terms of $F$-measure than those obtained by US or SW. Although the $F$-measure is improved by increasing the size of the majority class while keeping the size of the minority class fixed in UB, the performance of US does not depend on the size of the excess examples of the majority class. These are clearly different from the case of the naive bagging in training GLMs without considering the structure of class imbalance. That is, the ensembling and the regularization are different when the data structure is taken into account. The performance of SW is even worse, i.e., its performance on the $F$-measure degrades as the size of the majority class increases, leading to a catastrophic performance degradation when training from data with small minority class.
  • Furthermore, it is observed that UB is robust to the existence of the interpolation phase transition where data from linearly separable to inseparable training data, although the performance of the standard interpolator obtained by a single realization of training data is affected by such a phase transition as reported in (Mignacco et al., 2020).
@@ 1.1 Further related works
Learning under class imbalance is a classic and notorious problem in data mining and ML. In fact, the development of dealing with imbalanced data was selected as “10 CHALLENGING PROBLEMS IN DATA MINING RESEARCH” (YANG & WU, 2006) about 20 years ago in data mining community. Due to its practical relevance, plethora of methodological research has been continuously conducted and are summarized in reviews (He & Garcia, 2009; Longadge & Dongre, 2013; Johnson & Khoshgoftaar, 2019).
@@
Learning methods for imbalanced data can broadly be classified into two categories: cost- and data-level ones. In cost-sensitive methods, the cost function itself is modified to assign different costs to each class. A classical method includes weighting method that assigns a different constant multiplier for each data point depending on the class it belongs to. However, in the training of overparametrized models, it has been pointed out that a classical weighting methods fails and a more delicate handling of loss design is required to properly adjust the margins between classes (Lin et al., 2017; Khan et al., 2018; Byrd & Lipton, 2019; Ye et al., 2020; Kini et al., 2021). In data-level methods, the training data itself is modified to create a class-balanced dataset using resampling methods. The major approaches include US, which randomly discard a portion of the majority class data, and SMOTE (Chawla et al., 2002), which creates synthetic minority data points. Although the literature (Wallace et al., 2011) states that a combination of US and bagging is effective in learning overparametrized models, the comparison of other regularization techniques are not discussed.
@@
From a technical viewpoint, our work is an analysis of the sharp asymptotics of resampling methods in which the input dimension and the size of the dataset diverge at the same rate. Analyzing such a sharp asymptotics of ML methods in this asymptotic regime has been actively conducted since the 1990s, using techniques from statistical mechanics (Opper & Kinzel, 1996; Opper, 2001; Engel & Van den Broeck, 2001), and it has recently become one of the standard analytical approaches. The technical tools for analyzing such asymptotics include the replica method (Mézard et al., 1987; Gerace et al., 2020; Charbonneau et al., 2023; Montanari & Sen, 2024), convex Gaussian minimax theorem (Thrampoulidis et al., 2018), random matrix theory (Mei & Montanari, 2022), adaptive interpolation method (Barbier & Macris, 2019; Barbier et al., 2019), approximate message passing (Sur & Candès, 2019; Feng et al., 2022), and so on.
@@
Resampling methods have also been studied in the same proportional asymptotics. (Sollich & Krogh, 1995; Krogh & Sollich, 1997; Malzahn & Opper, 2001; 2002; 2003; LeJeune et al., 2020; Ando & Komaki, 2023; Bellec & Koriyama, 2024; Clarté et al., 2024; Patil & LeJeune, 2024) analyzes the performance of ensemble average of the linear models, and (Takahashi, 2023) analyzes the distribution of the ensemble average of the de-biased elastic-net estimator (Javanmard & Montanari, 2014b; a). However, these analyses do not address the structure of class imbalance. Ensembling in random feature models as a method to reduce the effect of random initialization are also considered in (D’Ascoli et al., 2020; Loureiro et al., 2022). Furthermore, the literature (Obuchi & Kabashima, 2019; Takahashi & Kabashima, 2019; 2020) derives the properties of the estimators trained on the resampled dataset in the analyses of the dynamics of approximate message passing algorithms that yields exact result in Gaussian or rotation-invariant feature data.
@@ 2 Problem setup
This section presents the problem setup and our interest in this work. The assumptions on the data generation process are first described, and then the estimator of the interest is introduced.
Let $D^{+}=\{(\mathbf{\boldsymbol{x}}_{\mu}^{+},1)\}_{\mu=1}^{M^{+}},\mathbf{\boldsymbol{x}}_{\mu}^{+}\in\mathbb{R}^{N}$ and $D^{-}=\{(\mathbf{\boldsymbol{x}}_{\nu}^{-},-1)\}_{\nu=1}^{M^{-}},\mathbf{\boldsymbol{x}}_{\nu}^{-}\in\mathbb{R}^{N}$ be the sets of positive and negative examples; in total we have $M=M^{+}+M^{-}$ examples. In this study, the positive class is assumed to be the minority class, that is, $M^{+}<M^{-}$. Also, it is assumed that positive and negative samples are drawn from a mixture models whose centroids are located at ±$\pm\mathbf{\boldsymbol{v}}/\sqrt{N}$ with $\mathbf{\boldsymbol{v}}\in\mathbb{R}^{N}$ being a fixed vector as follows:
@@
where the factor $1/\sqrt{N}$ is introduced to ensure that the logit $\mathbf{\boldsymbol{x}}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+B$ should be of ${\mathcal{O}}(1)$, and $f$ is a some nonlinear function. In the following, we denote $\mathbf{\boldsymbol{\theta}}=(\mathbf{\boldsymbol{w}},B)$ for the shorthand notation of the linear model’s parameter.
@@
$+\sum_{\nu=1}^{M^{-}}c_{\nu}^{-}l^{-}(\mathbf{\boldsymbol{x}}_{\mu}^{-}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+b)+\lambda\sum_{i=1}^{N}r(w_{i}),$ (4)
@@
$+\sum_{\nu=1}^{M^{-}}c_{\nu}^{-}l^{-}(\mathbf{\boldsymbol{x}}_{\mu}^{-}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+\bar{B})+\lambda\sum_{i=1}^{N}r(w_{i}),$ (5)
when the bias is fixed as $\bar{B}$ during the training. Here, $l^{+}$ and $l^{-}$ are the convex loss for positive and negative examples, respectively. For example, the choice of $l^{+}(x)=-\log(\sigma(x))$ and $l^{-}(x)=-\log(1-\sigma(x))$, with $\sigma(\cdot)$ being the sigmoid function, yields the cross-entropy loss. Also, $c_{\mu}^{+}\sim_{\rm iid}p_{c}^{+}$ and $c_{\nu}^{-}\sim_{\rm iid}p_{c}^{-}$ are the random coefficients that represent the effect of the resampling or the weighting. In the following we use the notation $\mathbf{\boldsymbol{c}}\equiv\{c_{\mu}^{+}\}_{\mu=1}^{M^{+}}\cup\{c_{\nu}^{-}\}_{\nu=1}^{M^{-}}$. Finally, $\lambda r(\cdot)$ is the convex regularizer for the weight vector with a regularization parameter $\lambda\in(0,\infty)$. By appropriately specifying the distribution $p_{c}^{\pm}$, the estimators (4) and (5) can represent estimators for different types of costs for resampled or reweighted data as follows:
@@
    where $\mu_{\ast}\in(0,1)$ be the resampling rate. $\mu_{\ast}=M^{+}/M^{-}$ is the natural choice to balance the size of the sizes of positive and negative examples.
@@
    $p_{c}^{+}(c^{-})$ $=\delta_{\rm d}(c^{+}-1),$ (8)
    $p_{c}^{-}$ $={\rm Poisson}(c^{-};\mu_{\ast}),$ (9)
    where $\mu_{\ast}\in(0,1)$ be the resampling rate.
@@
The ensemble average of the prediction for the input $\mathbf{\boldsymbol{x}}$ is thus obtained as follows:
@@
In this study, we focus on the case of subsampling as a method of resampling, since the the computational cost is usually smaller than the bootstrap.
The aim of this study is twofold: one is to characterize how the estimators (4) and (5) statistically depend on $D$ and $\mathbf{\boldsymbol{c}}$, and the other is to clarify the difference in performance with the UB obtained by (12) and the US or SW obtained by (4) or (5) with fixed $\mathbf{\boldsymbol{c}}$. We are interested in the generalization performance for both positive and negative inputs. Therefore, we will evaluate the generalization performance using the $F$-measure, the weighted harmonic mean of specificity and recall.
To characterize the properties of the estimators (4) and (5), we consider the large system limit where $N,M^{+},M^{-}\to\infty$, keeping their ratios as $(M^{+}/N,M^{-}/N)=(\alpha^{+},\alpha^{-})\in(0,\infty)^{2}$. In this asymptotic limit, the behavior of the estimators (4) and (5) can be sharply characterized as shown in the next section. We term this asymptotic limit as large system limit (LSL). Hereafter, $N\to\infty$ represents LSL as a shorthand notation to avoid cumbersome notation.

### v2

@@ Abstract
Under-bagging (UB), which combines under sampling and bagging, is a popular ensemble learning method for training classifiers on an imbalanced data. Using bagging to reduce the increased variance caused by the reduction in sample size due to under sampling is a natural approach. However, it has recently been pointed out that in generalized linear models, naive bagging, which does not consider the class imbalance structure, and ridge regularization can produce the same results. Therefore, it is not obvious whether it is better to use UB, which requires an increased computational cost proportional to the number of under-sampled data sets, when training linear models. Given such a situation, in this study, we heuristically derive a sharp asymptotics of UB and use it to compare with several other standard methods for learning from imbalanced data, in the scenario where a linear classifier is trained from a two-component mixture data. The methods compared include the under-sampling (US) method, which trains a model using a single realization of the subsampled data, and the simple weighting (SW) method, which trains a model with a weighted loss on the entire data. It is shown that the performance of UB is improved by increasing the size of the majority class while keeping the size of the minority fixed, even though the class imbalance can be large, especially when the size of the minority class is small. This is in contrast to US, whose performance does not change as the size of the majority class increases, and SW, whose performance decreases as the imbalance increases. These results are different from the case of the naive bagging when training generalized linear models without considering the structure of the class imbalance, indicating the intrinsic difference between the ensembling and the direct regularization on the parameters.
@@ 1 Introduction
In the context of binary classification, class imbalance refers to the situations where one class overwhelms the other in terms of the number of data points. Learning on class imbalanced data often occurs in real-world classification tasks, such as medical data analysis (Cohen et al., 2006), image classification (Khan et al., 2018), and so on. The primary goal when training classifiers on such imbalanced data is to achieve good generalization to both minority and majority classes. However, standard supervised learning methods without any special care often results in poor generalization performance for the minority class. Therefore, special treatments are necessary and have been one of the major research topics in the machine learning (ML) community for decades.
@@
Under-bagging (UB) (Wallace et al., 2011) is a popular and efficient method for dealing with a class imbalance that combines under-sampling (US) and bagging (Breiman, 1996). The basic idea of UB is to deal with class imbalance using under-sampling that randomly discards a portion of the majority class data to ensure an equal number of data points between the minority and the majority classes. However, this data reduction increases the variance of the estimator, because the total number of the training data is reduced. To mitigate this, UB aggregates estimators for different under-sampled data sets, similar to the aggregation performed in bootstrap aggregation (Breiman, 1996). US corresponds to a method of training a classifier on a single undersampled data. (Wallace et al., 2011) argues that UB is particularly effective for training of overparameterized models, where the number of the model parameters far exceeds the size of the training data.
@@
There is another line of approach, that is, cost-sensitive methods, which modify the training loss at each data point depending on the class it belongs to. However, it has recently been pointed out that training overparameterized models using cost-sensitive methods requires delicate design of loss functions (Lin et al., 2017; Khan et al., 2018; Byrd & Lipton, 2019; Ye et al., 2020; Kini et al., 2021) instead of the naive weighted cross-entropy (King & Zeng, 2001) and so on. In contrast to such delicate cost-sensitive approaches, UB has the advantage of being relatively straightforward to use, since it achieves complete class balance in any under-sampled dataset.
@@
Bagging is a natural approach to reduce the increased variance of weak learners, resulting from the smaller data size due to the resampling. However, unlike decision trees, which are the primary models used in bagging, neural networks can use simple regularization methods such as the ridge regularization to reduce the variance. In fact, it has recently been reported experimentally that naive bagging, which ignores the structure of the class balance, does not improve the classification performance of deep neural networks (Nixon et al., 2020). Furthermore, in generalized linear models (GLMs), it has been reported both theoretically (Sollich & Krogh, 1995; Krogh & Sollich, 1997; LeJeune et al., 2020; Du et al., 2023) and experimentally (Yao et al., 2021) that bagging with a specific sampling rate is exactly equivalent to the ridge regularization with a properly adjusted regularization parameter. These results indicate that naive bagging plays the role of an implicit regularization in linear models. Given these results, the question arises whether it is better to use UB, which requires an increased computational cost proportional to the number of under-sampled data sets, when training linear models or neural networks, rather than just using US and the ridge regularization.
@@
Given the above situation, this study aims to clarify whether bagging or ridge regularization should be used when learning a linear classifier from a class-imbalanced data. Specifically, we compare the performance of classifiers obtained by UB with those obtained by (i) the under-sampling (US) method, which trains a model using a single realization of the subsampled dataset, and (ii) the simple weighting (SW) method, which trains a model with a weighted loss on the entire data, when training a linear model on two-component mixture data, through the $F$-measure. To this end, we heuristically derive a sharp characterization of the classifiers obtained by minimizing the randomly reweighted empirical risk in the asymptotic limit where the input dimension and the parameter dimension diverge at the same rate, and then use it to analyze the performance of UB, US, and SW.
@@
  • A sharp characterization of the statistical behavior of the linear classifier is derived, where the classifier is obtained by minimizing the reweighted empirical risk function on two-component mixture data, in the asymptotic limit where input dimension and data size diverge proportionally (Claim 3). This includes the properties of the estimator obtained from the under-sampled data and the simple weighted loss function. There, the statistical property of the logits for a new input is described by a small finite set of scalar quantities that is determined as a solution of nonlinear equations, which we call self-consistent equations (Definition 1). The sharp characterization of the $F$-measure is also obtained (Claim 4). The derivation is based on the replica method of statistical physics (Mézard et al., 1987; Charbonneau et al., 2023; Montanari & Sen, 2024).
  • Using the derived sharp asymptotics, it is shown that the estimators obtained by UB have higher generalization performance in terms of $F$-measure than those obtained by US or SW. Although the performance of UB in terms of the $F$-measure is improved by increasing the size of the majority class while keeping the size of the minority class fixed, the performance of US does not depend on the size of the excess examples of the majority class (Section 4.1). This result is clearly different from the case of the naive bagging in training GLMs without considering the structure of the class imbalance, where the ensembling and ridge regularization yield the same performance. That is, the ensembling and the regularization are different when the data structure is taken into account. The performance of SW is even worse, i.e., its performance on the $F$-measure decreases as the size of the majority class increases, leading to a catastrophic performance degradation when training from data with a small minority class (Section 4.2).
  • Furthermore, it is shown that UB is robust to the existence of the interpolation phase transition, where data goes from linearly separable to inseparable training data (Section 4.1.1). In contrast, the performance of the interpolator obtained from a single realization of balanced data is affected by such a phase transition as reported in the previous study (Mignacco et al., 2020).
@@ 1.1 Further related works
Learning under class imbalance is a classic and notorious problem in data mining and ML. In fact, the development of dealing with imbalanced data was selected as “10 CHALLENGING PROBLEMS IN DATA MINING RESEARCH” (Yang & Wu, 2006) about 20 years ago in the data mining community. Due to its practical relevance, a lot of methodological research has been continuously conducted and are summarized in reviews (He & Garcia, 2009; Longadge & Dongre, 2013; Johnson & Khoshgoftaar, 2019; Mohammed et al., 2020).
@@
Learning methods for imbalanced data can broadly be classified into two categories: cost- and data-level ones. In cost-sensitive methods, the cost function itself is modified to assign different costs to each class. A classic method includes a weighting method that assigns a different constant multiplier to the loss at each data point depending on the class it belongs to. However, when training overparameterized models, it has been pointed out that a classic weighting method fails and a more delicate design of the loss function is required to properly adjust the margins between classes (Lin et al., 2017; Khan et al., 2018; Byrd & Lipton, 2019; Ye et al., 2020; Kini et al., 2021). In data-level methods, the training data itself is modified to create a class-balanced dataset using resampling methods. The major approaches include US (Drummond et al., 2003), which randomly discards some of the majority class data, and SMOTE (Chawla et al., 2002), which creates synthetic minority data points. Although the literature (Wallace et al., 2011) argues that a combination of US and bagging is effective for learning overparameterized models, the comparison with other regularization techniques is not discussed.
@@
Technically, our work is an analysis of the sharp asymptotics of resampling methods in which the input dimension and the size of the data set diverge at the same rate. The analysis of such sharp asymptotics of ML methods in this asymptotic regime has been actively carried out since the 1990s, using techniques from statistical physics (Opper & Kinzel, 1996; Opper, 2001; Engel & Van den Broeck, 2001). Recently, it has become one of the standard analytical approaches, and various mathematical techniques have also been developed to deal with it. The salient feature of this proportional asymptotic regime is that the macroscopic properties of the learning results, such as the predictive distribution and the generalization error, do not depend on the details of the realization of the training data111 In contrast, microscopic quantities such as each element of the weight vector fluctuate depending on each realization of the training data. , when using convex loss. That is, the fluctuations of these quantities with respect to the training data vanish, allowing us to make sharp theoretical predictions (Barbier et al., 2019; Charbonneau et al., 2023). The technical tools for analyzing such asymptotics include the replica method (Gerace et al., 2020; Charbonneau et al., 2023; Montanari & Sen, 2024), convex Gaussian minimax theorem (Thrampoulidis et al., 2018), random matrix theory (Mei & Montanari, 2022), adaptive interpolation method (Barbier & Macris, 2019; Barbier et al., 2019), approximate message passing (Sur & Candès, 2019; Feng et al., 2022), and so on.
@@
Resampling methods have also been studied in the same proportional asymptotics. (Sollich & Krogh, 1995; Krogh & Sollich, 1997; Malzahn & Opper, 2001; 2002; 2003; LeJeune et al., 2020; Ando & Komaki, 2023; Bellec & Koriyama, 2024; Clarté et al., 2024; Patil & LeJeune, 2023) analyze the performance of the ensemble average of the linear models, and (Takahashi, 2023) analyzes the distribution of the ensemble average of the de-biased elastic-net estimator (Javanmard & Montanari, 2014a; b). However, these analyses do not address the structure of the class imbalance. Ensembling in random feature models as a method to reduce the effect of random initialization are also considered in (D’Ascoli et al., 2020; Loureiro et al., 2022). Furthermore, the literature (Obuchi & Kabashima, 2019; Takahashi & Kabashima, 2019; 2020) derives the properties of the estimators trained on the resampled data set in the analysis of the dynamics of approximate message passing algorithms that yield exact results in Gaussian or rotation-invariant feature data.
@@ 2 Problem setup
This section presents the problem formulation and our interest in this work. First, the assumptions about the data generation process are described, and then the estimator of the interest is introduced.
Let $D^{+}=\{(\mathbf{\boldsymbol{x}}_{\mu}^{+},1)\}_{\mu=1}^{M^{+}},\mathbf{\boldsymbol{x}}_{\mu}^{+}\in\mathbb{R}^{N}$ and $D^{-}=\{(\mathbf{\boldsymbol{x}}_{\nu}^{-},-1)\}_{\nu=1}^{M^{-}},\mathbf{\boldsymbol{x}}_{\nu}^{-}\in\mathbb{R}^{N}$ be the sets of positive and negative examples; in total we have $M=M^{+}+M^{-}$ examples. In this study, the positive class is assumed to be the minority class, that is, $M^{+}<M^{-}$. Also, it is assumed that positive and negative samples are drawn from a mixture model whose centroids are located at ±$\pm\mathbf{\boldsymbol{v}}/\sqrt{N}$ with $\mathbf{\boldsymbol{v}}\in\mathbb{R}^{N}$ being a fixed vector as follows:
@@
where the factor $1/\sqrt{N}$ is introduced to ensure that the logit $\mathbf{\boldsymbol{x}}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+B$ should be of ${\mathcal{O}}(1)$, and $f$ is a some nonlinear function. In the following, we denote $\mathbf{\boldsymbol{\theta}}$ for the shorthand notation of the linear model’s parameter.
@@
$+\sum_{\nu=1}^{M^{-}}c_{\nu}^{-}l^{-}(\mathbf{\boldsymbol{x}}_{\nu}^{-}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+b)+\lambda\sum_{i=1}^{N}r(w_{i}),$ (4)
@@
$+\sum_{\nu=1}^{M^{-}}c_{\nu}^{-}l^{-}(\mathbf{\boldsymbol{x}}_{\nu}^{-}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+\bar{B})+\lambda\sum_{i=1}^{N}r(w_{i}),$ (5)
when the bias is fixed as $\bar{B}$ during the training. Here, $l^{+}$ and $l^{-}$ are the convex loss for positive and negative examples, respectively. For example, the choice of $l^{+}(x)=-\log(\sigma(x))$ and $l^{-}(x)=-\log(1-\sigma(x))$, with $\sigma(\cdot)$ being the sigmoid function, yields the cross-entropy loss. Also, $c_{\mu}^{+}\sim_{\rm iid}p_{c}^{+}$ and $c_{\nu}^{-}\sim_{\rm iid}p_{c}^{-}$ are the random coefficients that represent the effect of the resampling or the weighting. In the following, we use the notation $\mathbf{\boldsymbol{c}}\equiv\{c_{\mu}^{+}\}_{\mu=1}^{M^{+}}\cup\{c_{\nu}^{-}\}_{\nu=1}^{M^{-}}\sim p_{c}$. Finally, $\lambda r(x)=\lambda x^{2}/2$ is the ridge regularization for the weight vector with a regularization parameter $\lambda\in(0,\infty)$.
By appropriately specifying the distribution $p_{c}^{\pm}$, the estimators (4) and (5) can represent estimators for different types of costs for resampled or reweighted data as follows:
@@
    where $\mu_{\ast}\in(0,1]$ be the resampling rate. $\mu_{\ast}=M^{+}/M^{-}$ is the natural choice to balance the size of the sizes of positive and negative examples.
@@
    $p_{c}^{+}(c^{+})$ $=\delta_{\rm d}(c^{+}-1),$ (8)
    $p_{c}^{-}$ $={\rm Poisson}(\mu_{\ast}),$ (9)
    where $\mu_{\ast}\in(0,1]$ be the resampling rate.
@@
The ensemble average of the prediction for a new input $\mathbf{\boldsymbol{x}}$ is thus obtained as follows:
@@
In this study, we focus on the case of subsampling as a method of resampling (6)-(7), since the the computational cost is usually smaller than the bootstrap.
The purpose of this study is twofold: the first is to characterize how the prediction $\mathbf{\boldsymbol{x}}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)/\sqrt{N}+\hat{B}(\mathbf{\boldsymbol{c}},D)~{}({\rm or~{}}\bar{B})$ depends statistically on $\mathbf{\boldsymbol{x}}$ and $\mathbf{\boldsymbol{c}}$, and the other is to clarify the difference in performance with the UB obtained by (12) and the US or SW obtained by (4) or (5) with fixed $\mathbf{\boldsymbol{c}}$. We are interested in the generalization performance for both positive and negative inputs. Therefore, we will evaluate the generalization performance using the $F$-measure, the weighted harmonic mean of specificity and recall.
To characterize the properties of the estimators (4) and (5), we consider the large system limit where $N,M^{+},M^{-}\to\infty$, keeping their ratios as $(M^{+}/N,M^{-}/N)=(\alpha^{+},\alpha^{-})\in(0,\infty)^{2}$. In this asymptotic limit, the behavior of the estimators (4) and (5) can be sharply characterized as shown in the next section. We call this asymptotic limit as large system limit (LSL). In the following, $N\to\infty$ represents the LSL as a shorthand notation to avoid cumbersome notation.